# A Closed-Form Solution for Kernel Adaptive Filtering

### v2

@@ 3.2 Covariance kernel for the approximate MMSE Solution with the explicit feature map
The solution in section 2.3 requires the definition of a family of r.v.s. with finite second-order moments. Our truncated approximation that computes an explicit feature map provide us with an alternative family of r.v.s $\{\phi_{m}(\bm{X}^{D}_{t}),(m,t)\in\mathbb{N}\times T\}$, abbreviated as $\phi(\bm{X}^{D})$ as our family of functions with finite second-order moments. Next, we will show how to compute the covariance kernel, $\bm{U}(s,t,m,n)$, and define the inner product in $\mathcal{H}_{\bm{U}}$. Then we will show how to compute the projection of some desired r.v. into $\mathcal{H}_{\bm{U}}$ giving the nonlinear MMSE solution.
@@
$\bm{U}(t,s,m,n)=\mathbb{E}[\phi_{m}(\bm{X}_{t}^{D})\phi_{n}(\bm{X}_{s}^{D})]\quad t,s\in T\quad m,n\in[1,M]$ (24)
Then, we can approximate random variables in $L^{2}(\phi(\bm{X}^{D}))$ as defined in (19), as follows:
$\begin{split}W\approx\sum_{q=1}^{m_{W}}\sum\limits_{m=1}^{M}\sum_{i=1}^{n_{W}}&a_{iq}\phi_{m}(x_{i})\phi_{m}(\bm{X}_{t_{q}}^{D})=\sum_{q=1}^{m_{W}}\sum\limits_{m=1}^{M}A_{q,m}\phi_{m}(\bm{X}_{t_{q}}^{D})\\ &A_{q,m}=\sum_{i=1}^{n_{W}}a_{iq}\phi_{m}(x_{i})\end{split}$ (25)
With inner product given by:
$\begin{split}\langle W,V\rangle_{L^{2}}=\mathbb{E}[WV]&=\sum_{q=1}^{m_{W}}\sum\limits_{m=1}^{M}\sum\limits_{p=1}^{m_{V}}\sum\limits_{n=1}^{M}A_{q,m}B_{p,n}\mathbb{E}[\phi_{m}(X_{t_{q}})\phi_{n}(X_{t_{p}})]\\ &=\sum_{q=1}^{m_{W}}\sum\limits_{m=1}^{M}\sum\limits_{p=1}^{m_{V}}\sum\limits_{n=1}^{M}A_{q,m}B_{p,n}\bm{U}(t_{q},t_{p},m,n)\end{split}$ (26)
@@
Finally, we can define the projection of a desired r.v. $Z$ into $\mathcal{H}_{\bm{U}}$ using the crosscovariance function.
$\tilde{\rho}(t,m)=\mathbb{E}[Z\phi_{m}(\bm{X}^{D}_{t})]$ (27)
The feature space representation of this function can be given as,
$\bm{\rho}=\mathbb{E}[Z\phi(\bm{X}_{t}^{DL})]\quad\phi(\bm{X}^{DL}_{t}):=\begin{bmatrix}\phi(\bm{X}^{D}_{t})\\ \vdots\\ \phi(\bm{X}^{D}_{t-(L-1)})\end{bmatrix}\in\mathbb{R}^{M\cdot L}.$ (28)
where
@@
$\hat{Z}=\phi(\bm{X}_{t}^{DL})^{\top}\mathbf{U}^{\dagger}\bm{\rho}$ (29)
Notice that this equation has the same form as the linear MMSE solution given in equation (10) except that the number of dimensions is larger for the nonlinear case. While the number of dimensions in the linear MMSE solution is related only to time, the number of dimensions in the nonlinear solution is related to both time and feature space index as a result of the RKHS we employ (see Table 1 for a comparison between all spaces introduced here).
@@
Table 1: Brief description of the different RKHSs used so far.
@@ 3.3 Error Analysis
Now we will discuss some sources of error in the FWF solution. First, it is essential to remember that we can only calculate the FWF solution in a finite-dimensional RKHS, therefore $\mathcal{H}_{S}$, and transitively $\mathcal{H}_{\bm{U}}$, can only be universal if $K$ approaches infinity and $D$ approaches the true model order. This deficiency of $\mathcal{H}_{\bm{U}}$ can be a source of error. In [3], error bounds for the approximation made by $S(\cdot,\cdot)$ to the universal Gaussian kernel are given. Second, we still need to select the Gaussian kernel size, which in the FWF only affects the precision of the results for finite training datasets. Third, since we are quantifying the joint distribution between pairs of samples projected into an RKHS and taking their mean, implicitly, we are assuming that the system we are modeling is strictly stationary. Any deviation from this stationary assumption can cause error. Fourth, there can be numerical error associated with the conditioning of the $\mathbf{U}$ matrix.
In [25] a theoretical MMSE for the solution in $\mathcal{H}_{R}$ is given with the equation
$\mathbb{E}[|Z-\mathbb{E}[Z|L^{2}(X)]|^{2}]=\mathbb{E}[|Z|^{2}]-\langle\rho,\rho\rangle_{\mathcal{H}_{R}}.$ (30)
We can extend this theoretical MMSE to $\mathcal{H}_{\bm{U}}$ by measuring the distance between the projection of $Z$ into $\mathcal{H}_{\bm{U}}$ and $Z$ in a similar manner to equation (30).
@@
This equation gives the MMSE for the solution in the finite-dimensional RKHS, $\mathcal{H}_{\bm{U}}$. Since the projection of $Z$ in $\mathcal{H}_{\bm{U}}$ is orthogonal, the difference between the power of the desired response and the norm squared of its projection in $\mathcal{H}_{\bm{U}}$ is the expected MSE. Therefore, it is quite easy to select the hyper parameters of the modeling ($M$, $L$ and kernel size) to achieve the desired error. This is markedly different from KAF and other filtering methods, where we have to evaluate the MSE on the training set by directly comparing model outputs with desired responses for different hyperparameters. Therefore, discounting numerical error, we can say that all the sources of error mentioned above are combined in the inner product calculation of the space $\mathcal{H}_{\bm{U}}$.
@@
Figure 2 shows the theoretical MMSE, the absolute difference between theoretical and calculated MSE in the training set, and the test set MSE as a function of the FWF hyperparameters for prediction on the nonlinear chaotic Mackey-Glass time series with a kernel size of 0.25. As can be expected the larger the number of lags and dimension of the explicit RKHS, the smaller the error. This figure also shows that the training set MSE follows the theoretical error almost exactly for a large range of hyperparameter values. Figure 2 also shows the test set MSE as a function of $M$ and $L$. We can observe that the test set performance largely follows the training MSE, but for large delays and number of eigenfunctions flattens, which is explained by over-fitting. Since the FWF is a linear model, techniques for model order estimation [35] can be applied to improve generalization, but we leave this to future work.
@@
Figure 2: ||Training - Theoretical|| MSE in log scale (middle), Theoretical MSE (left), and Test MSE (right) as a function of $K$ and $L$ for prediction of Mackey-Glass time series.
3.4 The Functional Wiener Equation
Thus far, we have given the solution for a nonlinear MMSE filter in $\mathcal{H}_{U}$ by extending Parzen’s ideas. It is interesting to see the ramifications of this formulation on the system of equations solved by the linear Wiener Filter. To this end, we will now develop an analog to this system of equations, which describes the problem we implicitly solve with the Functional Wiener Filter.
In this discussion, we will assume that the kernel-based integral operators in question are Hilbert-Schmidt. We start by examining the operator related to the kernel $U$. An operator over a Hilbert space can be represented as an integral operator defined by its corresponding kernel function. This representation is coordinate-free. Alternatively, a matrix with dimensions equal to the number of dimensions in the Hilbert Space can be used, here the values in this matrix depend on some chosen coordinate system. In this work, the $\mathbf{U}$ matrix is a matrix representation of an operator with respect to the coordinate system(feature space) of ${\mathcal{H}_{S}}^{L}$. Using the relationship between $\mathbf{U}$ and $\bm{U}(s,t,m,n)$ given in Appendix A we can recover an operator over $\mathcal{H}_{S}$ for each pair of time indices $s,t$
$U(s,t,x,y)\approx\sum_{n=1}^{M}\sum_{m=1}^{M}\bm{U}(t,s,n,m)\phi_{n}(x)\phi_{m}(y)$ (32)
where $x,y\in\mathcal{X}\subset\mathbb{R}^{D}$, for specific values of $s,t\in T$ and $n,m\in[1,M]$. We can think of $U(s,t,x,y)$ as a collection of operators over functions of the RKHS $\mathcal{H}_{\kappa}$ induced by the kernel $\kappa$, indexed with pairs of indices $(s,t)$. Alternatively, when we only consider $L$ relative times, we could view $U(s,t,x,y)$ as a single operator over $\mathcal{H}_{\kappa}^{L}$. The equation implicitly solved in (29) can then be given as
$\sum_{t=0}^{L-1}\left[\int_{\mathcal{X}}w(t,x)U(s,t,x,y)dx\right]=\rho(s,y)$ (33)
where $w,\rho\in{\mathcal{H}_{\kappa}}^{L}$. Note that $\rho(s,y)$ is the crosscovariance function defined in (15).
When we approximate $U$ based on the truncated kernel $S$ which is represented by the matrix $\mathbf{U}$ and multiply this by the vector $\bm{\rho}$ we solve for $\mathbf{w}^{*}=\mathbf{U}^{\dagger}\bm{\rho}$, which is the vector approximation of the function, $w^{*}(t,x)$ that solves (33). Equation (33) is a direct nonlinear analog to the system of equations solved by the linear Wiener Filter in discrete time. For this reason, we call it the Functional Wiener Filter equation.
In principle, the extension of equation (33) for continuous time can be written as,
$\int_{a}^{b}\left[\int_{\mathcal{X}}w(t,x)U(s,t,x,y)dx\right]dt=\rho(s,y)$ (34)
where $a\leq s\leq b$. However, the techniques detailed in section 3.2 do not extend to the continuous time operator.
@@ 3.5 The Functional Wiener Filter as a Difference Equation
Similar to the linear Wiener Filter, the Functional Wiener Filter can be interpreted as the optimal solution under the assumption of a difference equation with a specific form. In the linear case this difference equation corresponds to an FIR linear filter, with the form
$\hat{z_{t}}=\sum^{L-1}_{\tau=0}w_{\tau}x_{t-\tau}\quad w_{\tau},x_{t-\tau}\in\mathbb{R}$ (35)
@@
In the case of the Functional Wiener Filter we replace scalar multiplication with nonlinear functions. Due to sample embeddings, these nonlinear functions may be defined over multiple samples rather than single samples.
$\hat{z_{t}}=\sum^{L-1}_{\tau=0}f_{\tau}(\mathbf{x}_{t-\tau}^{D})\quad f_{\tau}\in\mathcal{H}_{S},\mathbf{x}_{t}^{D}\in\mathbb{R}^{D}$ (36)
@@
The easiest way to explain how the FWF defines a difference equation is by first calculating,
$\mathbf{w}^{*}=\mathbf{U}^{\dagger}\bm{\rho}$ (37)
Then the FWF solution can be written as,
$Z_{t}=\phi(\bm{X}_{t}^{DL})^{\top}\mathbf{U}\bm{\rho}=\phi(\bm{X}^{DL}_{t})^{\top}\mathbf{w}^{*}$ (38)
The rightmost expression in equation 38 can be written as,
$\hat{z}_{t}=\phi(\mathbf{x}^{DL}_{t}))^{\top}\mathbf{w}^{*}=\sum_{\tau=0}^{L-1}\phi(\mathbf{x}^{D}_{t-\tau})^{\top}\mathbf{w}_{M\tau:(M\tau)+(M-1)}^{*}=\sum_{\tau=0}^{L-1}f^{*}_{\tau}(\mathbf{x}_{t}^{D})$ (39)
Therefore the subvectors $\mathbf{w}_{M\tau:(M\tau)+(M-1)}^{*}$ are the feature space representations of $f^{*}_{\tau}$.
@@
Note that $\mathcal{H}_{S}$ becomes universal on $\mathbb{R}^{D}$ when $K$ goes to infinity. Every solution found using the Functional Wiener Filter follows this form. This highlights a distinction between the FWF and other KAF methods found in [16], [15], [6], and [17]. The FWF is defined over $\mathbb{R}^{D}\times T$ (time and space) rather than just over $\mathbb{R}^{D}$. The interpretation of the FWF as a nonlinear difference equation gives a clearer picture of what the FWF learns and can provide insight into the dynamics of the model that generates the input time series and the failure states of the learned models. In the experimental section, we will demonstrate how this interpretation can be used to extract a difference equation from data. Note that this type of interpretation is not possible with any other KAF method mentioned above.
4 Experiments
@@
We will compare the FWF solution derived above in two important applications: nonlinear mapping of one time series into another (as required in system identification) and nonlinear time series prediction. The models used for comparison are the linear Wiener Filter (WF) [38], Gaussian Process Regression (GPR) [40], Kernel Least-Mean Squares (KLMS) [16], Extended Kernel Recursive Least-Squares (KRLS) [15], Kernel Ridge Regression (KRR) [10], and Augmented Space Linear Model (ASLM) [27], which is a linear model in the joint space of the data and desired response, improving slightly the performance of the Wiener solution. The error bars in figures 4,5,6 represent the range of test MSE(i.e. they show the worst test MSE and best test MSE).
@@
Table 2: Computational Complexity Comparison for both training and evaluation: $M=\binom{D+K}{K}$ (number of dimension in $\mathcal{H}_{S}$), N (number of training samples), L (window size)
Table 2 shows a comparison of the computational complexity between the different methods. Since the FWF utilizes the explicit finite-dimensional RKHS, the computation cost of evaluating the model is decoupled from the number of training samples used, as in the Wiener filter. The hyperparameters that affect the computational complexity of the FWF are $D$, $K$, and $L$, and if $N>>ML$, then the $\mathcal{O}(M^{2}L^{2}N)$ will dominate the training complexity. Note that the use of sample embeddings will increase the number of dimensions in $\mathcal{H}_{S}$; therefore, this generalization does not come without additional computational cost.
For each experiment, a grid search across the hyperparameters of each model was performed. However, for brevity, we give the finer details of the other searches in the supplementary materials.

### v3

@@ 3.2 Covariance kernel for the approximate MMSE Solution with the explicit feature map
The solution in section 2.2 requires the definition of a family of r.v.s. with finite second-order moments. Our truncated approximation that computes an explicit feature map provides us with an alternative family of r.v.s. $\{\phi_{m}(\bm{X}^{D}_{t}),(m,t)\in[1,M]\times T\}$, abbreviated as $\phi(\bm{X}^{D})$ as our family of functions with finite second-order moments. Now, we follow the steps detailed in section 2.2. First, we define the covariance kernel, $\bm{U}(s,t,m,n)$ and the inner product in $\mathcal{H}_{\bm{U}}$. Then we demonstrate how to compute the cross-covariance function, yielding the nonlinear MMSE solution in $\mathcal{H}_{\mathbf{U}}$.
@@
$\bm{U}(t,s,m,n)=\mathbb{E}[\phi_{m}(\bm{X}_{t}^{D})\phi_{n}(\bm{X}_{s}^{D})],\qquad t,s\in T\qquad m,n\in[1,M].$ (10)
We can represent any random variable $W\in L^{2}(\phi(\bm{X}^{D}))$ as,
$W=\sum_{q=1}^{m_{W}}\sum\limits_{m=1}^{M}\sum_{i=1}^{n_{W}}a_{iq}\phi_{m}(x_{i})\phi_{m}(\bm{X}_{t_{q}}^{D})=\sum_{q=1}^{m_{W}}\sum\limits_{m=1}^{M}A_{q,m}\phi_{m}(\bm{X}_{t_{q}}^{D}),$ (11)
where $A_{q,m}=\sum_{i=1}^{n_{W}}a_{iq}\phi_{m}(x_{i})$. The inner product is given by,
$\begin{split}\langle W,V\rangle_{L^{2}}=\mathbb{E}[WV]&=\sum_{q=1}^{m_{W}}\sum\limits_{m=1}^{M}\sum\limits_{p=1}^{m_{V}}\sum\limits_{n=1}^{M}A_{q,m}B_{p,n}\mathbb{E}[\phi_{m}(X_{t_{q}})\phi_{n}(X_{t_{p}})]\\ &=\sum_{q=1}^{m_{W}}\sum\limits_{m=1}^{M}\sum\limits_{p=1}^{m_{V}}\sum\limits_{n=1}^{M}A_{q,m}B_{p,n}\bm{U}(t_{q},t_{p},m,n).\end{split}$ (12)
@@
Finally, we can define the projection of a desired r.v. $Z$ into $\mathcal{H}_{\bm{U}}$ using the cross-covariance function, $\tilde{\rho}(t,m)=\mathbb{E}[Z\phi_{m}(\bm{X}^{D}_{t})]$. The feature space representation of this function can be given as,
$\bm{\rho}=\mathbb{E}[Z\phi(\bm{X}_{t}^{DL})],\qquad\phi(\bm{X}^{DL}_{t}):=\begin{bmatrix}\phi(\bm{X}^{D}_{t})\\ \vdots\\ \phi(\bm{X}^{D}_{t-(L-1)})\end{bmatrix}\in\mathbb{R}^{M\cdot L}.$ (13)
@@
$\hat{Z}=\langle\phi(\mathbf{X}_{t}),\tilde{\rho}\rangle_{\mathcal{H}_{\bm{U}}}=\phi(\bm{X}_{t}^{DL})^{\top}\mathbf{U}^{\dagger}\bm{\rho}.$ (14)
Notice that this equation has the same form as the linear MMSE solution given in equation (3) except that the number of dimensions is larger for the nonlinear case. While the number of dimensions in the linear MMSE solution is related only to time, the number of dimensions in the nonlinear solution is related to both time and dimensionality of $\phi(\cdot)$ as a result of the RKHS we employ. See Table 1 for a comparison between all RKHSs introduced thus far.
@@
Table 1: Brief description of the different RKHSs used so far. Note $\mathcal{H}_{\bm{U}}$ is in practice truncated.
@@ 3.3 Error Analysis
The FWF solution has several error sources. First, since calculations occur in a finite-dimensional RKHS, $\mathcal{H}_{S}$ and transitively $\mathcal{H}_{U}$ are only universal as K approaches infinity, D equals the true model order, and L equals one. Error bounds for $S(\cdot,\cdot)$’s approximation to the universal Gaussian kernel are given in [29]. Second, the proper selection of the kernel size remains necessary, affecting precision with finite training datasets. Third, by quantifying the joint distribution between sample pairs projected into the RKHS and taking their mean, we implicitly assume strict stationarity. Any deviations from this assumption may introduce error. Finally, numerical error can arise from the conditioning of the $\mathbf{U}$ matrix.
It may seem that all these approximations are challenging to quantify; however, the optimal solution provides a direct means for evaluating the MSE. In fact, since the optimal solution is an orthogonal projection, we can calculate the cumulative approximation MSE by simply measuring the distance between the projection of $Z$ into $\mathcal{H}_{\bm{U}}$ (i.e. $\tilde{\rho}$) and $Z$ itself.
@@
Since the projection of $Z$ in $\mathcal{H}_{\bm{U}}$ is orthogonal, the difference between the power of the desired response and the norm squared of its projection in $\mathcal{H}_{\bm{U}}$ is the expected MSE. Therefore, it is quite easy to select the hyperparameters of the model ($K$, $L$, $D$, and kernel size) to meet the minimum error specification. This is markedly different from KAF and other filtering methods, where we have to evaluate the MSE on the training set by directly comparing model outputs with desired responses for different hyperparameters. Discounting numerical error, all the sources of error mentioned above are combined in the inner product calculation of the space $\mathcal{H}_{\bm{U}}$.
@@
Figure 2 shows the theoretical MMSE, the absolute difference between theoretical and empirical MSE in the training set, and the test set MSE as a function of the FWF hyperparameters for prediction on the nonlinear chaotic Mackey-Glass time series introduced in 4.2. These empirical results confirm that the theoretical MMSE closely matches empirical error across a wide range of hyperparameters, and test set MSE largely follows training set MSE. Since the FWF is a linear model in the RKHS, techniques for model order estimation [32] can be applied to improve generalization, but we leave this to future work.
@@
Figure 2: ||Training - Theoretical|| MSE in log scale (middle), Theoretical MSE (left), and Test MSE (right) as a function of $M$ and $L$ for prediction of Mackey-Glass time series.
@@ 3.4 The Functional Wiener Filter as a Difference Equation
Similar to the linear Wiener Filter, the Functional Wiener Filter can be interpreted as the optimal solution under the assumption of a difference equation with a specific form. The difference equation assumed by the Functional Wiener Filter is similar to a linear FIR filter except we replace scalar multiplication with nonlinear functions. Due to sample embeddings, these nonlinear functions may be defined over multiple samples rather than single samples.
$\hat{z_{t}}=\sum^{L-1}_{\tau=0}f_{\tau}(\mathbf{x}_{t-\tau}^{D})\qquad f_{\tau}\in\mathcal{H}_{S},\mathbf{x}_{t}^{D}\in\mathbb{R}^{D}$ (16)
@@
The easiest way to see how the FWF defines a difference equation is by first calculating $\mathbf{w}^{*}=\mathbf{U}^{\dagger}\bm{\rho}$. Then the FWF solution can be written as,
$Z_{t}=\phi(\bm{X}_{t}^{DL})^{\top}\mathbf{U}^{\dagger}\bm{\rho}=\phi(\bm{X}^{DL}_{t})^{\top}\mathbf{w}^{*}.$ (17)
The expression $\phi(\bm{X}^{DL}_{t})^{\top}\mathbf{w}^{*}$ in equation 17 can be written as,
$\hat{z}_{t}=\sum_{\tau=0}^{L-1}\phi(\mathbf{x}^{D}_{t-\tau})^{\top}\mathbf{w}_{M\tau:(M\tau)+(M-1)}^{*}=\sum_{\tau=0}^{L-1}f^{*}_{\tau}(\mathbf{x}_{t-\tau}^{D}).$ (18)
@@
Therefore the subvectors, $\mathbf{w}_{M\tau:(M\tau)+(M-1)}^{*}$, are the feature space representations of $\{f^{*}_{\tau}\}_{\tau=0}^{L-1}$. Note that $\mathcal{H}_{S}$ becomes universal on $\mathbb{R}^{D}$ when $K$ goes to infinity. Every solution found using the Functional Wiener Filter follows this form. This highlights a distinction between the FWF and other KAF methods found in [3, 33, 4, 2]. The FWF is defined over $\mathbb{R}^{D}\times T$ (time and space) rather than just over $\mathbb{R}^{D}$. In the experimental section, we will demonstrate how this interpretation can be used to extract a difference equation from data. To our knowledge, this type of interpretation is not possible with any other KAF method.
3.5 Hyperparameters and Trade-offs
The hyperparameters of the FWF are sample embedding size ($D$), window size ($L$), truncation point ($K$), kernel size ($\sigma$), and regularization parameter (implicit in the Moore-Penrose pseudo inverse). The kernel size plays the same role as in other KAF methods. See [34] for a discussion on kernel size selection. The regularization parameter also plays a standard role, controlling the conditioning for the inversion of $\mathbf{U}$.
From a high level, the roles of $D$, $L$, and $K$ are similar; they increase model capacity and generalize $\mathcal{H}_{\mathbf{U}}$. The specific roles of $D$ and $L$ are best understood using (18). $D$ controls the domain of each function, $f^{*}_{\tau}$, while $L$ controls the memory depth of the system. $K$ affects model capacity by controlling the highest degree considered in our truncated Taylor series, which ultimately affects the hypothesis space for each $f_{\tau}^{*}$. If $D$, $L$, or $K$ are too small, then a satisfactory solution may not exist in $\mathcal{H}_{\bm{U}}$. Conversely, if they are too large, you unnecessarily increase the computational complexity, susceptibility to noise, and amount of data necessary for estimating the FWF well. Therefore, practitioners should aim to use the smallest values for $D$, $L$, and $K$ that permit the FWF to fit the data well.
4 Experiments and Simulations
@@
We now test the FWF solution in two important applications: nonlinear mapping of one time series into another (as required in system identification) and nonlinear time series prediction. The models used for comparison are the linear Wiener Filter (WF) [19], Gaussian Process Regression (GPR) [35], Kernel Least-Mean Squares (KLMS) [3], Extended Kernel Recursive Least-Squares (KRLS) [33], Kernel Ridge Regression (KRR) [36], and Augmented Space Linear Model (ASLM) [37]. The Gaussian kernel was used for all kernel methods.
For each experiment, a grid search was conducted across the hyperparameters of each model. The best results for each method are presented. For brevity, we give the finer details of the searches and final hyperparameter settings in the supplementary materials. An additional experiment on forecasting the Lorenz system is also included in the supplementary materials.
@@
Table 2: Computational Complexity Comparison for both training and evaluation: $M=\binom{D+K}{K}$ (number of dimensions in $\mathcal{H}_{S}$), N (number of training samples), L (window size)
Table 2 shows a comparison of the computational complexity between the different methods. The hyperparameters that affect the computational complexity of the FWF are $D$, $K$, and $L$. The FWF training complexity is the highest, proportional to the cube of the product of window size and dimension; however, the FWF is a batch method, making the computation in the training step parallelizable. In practice, it is often the case that $N>>ML$, so the $\mathcal{O}(M^{2}L^{2}N)$ will dominate the training complexity, which is similar to the linear case. While training complexity can be large, test time complexity is untethered from N, similar to the WF, which is a great computational advantage for low-power computation.